Provide step-by-step solutions to the given problem below. Write your answers on


the spaces provided.


Consider all samples of size 2 from the population 2, 5, 6, 9, 11, and 13. Compute


the mean, and variance of the sampling distribution of the sample mean.


First step:


Second step


a)


b)


Third step


Fourth step


Fifth step


Sixth step


a)


b)


Seventh step


a)


b)


c)


d)






1
Expert's answer
2022-03-29T07:38:16-0400



"The\,\,mean\,\,is\,\,\\\mu =\frac{23}{3}\\The\,\,variance\,\,is\\\sigma ^2=\frac{\sum{{\bar{x}_i}^2}}{15}-\mu ^2=\\=\frac{3.5^2+4^2+2\cdot 5.5^2+6.5^2+7^2+2\cdot 7.5^2+8^2+8.5^2+9^2+9.5^2+10^2+11^2+12^2}{15}-\left( \frac{23}{3} \right) ^2=\\=\frac{50}{9}"